# Leveraging Audio and Text Modalities in Mental Health: A Study of LLMs Performance

### v1

@@ Abstract
Mental health disorders such as depression and PTSD are increasingly prevalent worldwide, creating an urgent need for innovative tools to support early diagnosis and intervention. This study explores the potential of Large Language Models (LLMs) in multimodal mental health diagnostics, specifically for detecting depression and PTSD (Post Traumatic Stress Disorder) through text and audio modalities. Using the E-DAIC dataset, we compare text and audio modalities to investigate whether LLMs can perform equally well or better with audio inputs, assessing their effectiveness in capturing both vocal cues and linguistic patterns. We further examine the integration of both modalities to determine if this can enhance diagnostic accuracy, which generally results in improved performance metrics. Our analysis specifically utilizes custom-formulated metrics—Modal Superiority Score (MSS) and Disagreement Resolvement Score (DRS)—to evaluate how combined modalities influence model performance. The Gemini 1.5 Pro model achieves the highest scores in binary depression classification when using the combined modality, with an F1 score of 0.67 and a Balanced Accuracy (BA) of 77.4%, assessed across the full dataset. These results represent an increase of 3.1% over its performance with the text modality and 2.7% over the audio modality, highlighting the effectiveness of integrating modalities to enhance diagnostic accuracy. Similarly, with a BA of 77% and an F1 score of 0.68, the GPT-4o mini demonstrates significant success in classifying PTSD. Notably, all results are obtained in zero-shot inferring, highlighting the robustness of the models without requiring task-specific fine-tuning. To explore the impact of different configurations on model performance, we conduct binary, severity, and multiclass tasks using both zero-shot and few-shot prompts, examining the effects of prompt variations on performance. The results reveal that models such as Gemini 1.5 Pro in text and audio modalities, and GPT-4o mini in the text modality, often surpass other models in balanced accuracy and F1 scores across multiple tasks. This study highlights the promising role of LLMs in clinical settings for mental health assessment, emphasizing the need for advancements in LLM-based diagnostics for multimodal mental health applications.
@@ 1 Introduction
In recent years, models like BERT[2] have advanced rapidly, demonstrating significant capabilities in tasks ranging from natural language processing to decision-making in specialized domains." An Overview of Large Language Models (LLMs)" (2023)[3], this paper highlights the capabilities of LLMs as they are trained on vast and diverse corpora, enabling them to capture complex patterns in language. Their training allows them to generate coherent and contextually relevant text, making them highly adaptable across various applications. In healthcare, for example, LLMs can assist in diagnosing diseases, summarizing patient records, and even providing support for therapeutic interventions. Their ability to generalize across tasks has made them valuable in different fields, significantly enhancing efficiency and scalability.
LLMs show a remarkable ability to process and analyze vast amounts of text data, identifying and predicting psychiatric conditions by detecting patterns and subtle linguistic cues in patient communication. They excel at providing timely, scalable, and personalized assessments, aiding mental health professionals in diagnosis and treatment planning. LLMs provide scalable, efficient, and objective assessments, enhancing diagnostic accuracy and personalization of treatment. However, they must be used carefully, as they may generate inaccurate responses or lead to over-reliance, requiring continuous monitoring to ensure safety and effectiveness[4][5].
Recent LLM advancements have enabled models to process not only text but also audio inputs directly[6], providing valuable insights into mental health. Some models convert audio to text via speech recognition system [7], analyzing the linguistic content for symptoms of mental disorders. However, newer models can directly analyze audio by detecting differences in tone, cadence, and speech patterns, which can reflect emotional states more accurately than text alone [8]. This approach allows for deeper insights into a patient’s mental health, potentially leading to more precise and early diagnosis of conditions like depression or anxiety.
In this paper, we compare two approaches for mental illness detection: text modality and audio modality, using different models to analyze each. Models processing text evaluate written transcripts to identify linguistic patterns, while models processing audio analyze direct raw speech to capture features such as tone and cadence. Additionally, we explore the integration of both modalities to determine if this enhances model performance. To quantitatively assess how the combined modality either increases or decreases performance, we utilize our specially formulated metrics: the MSS and DRS. By evaluating both individual and integrated modalities, we aim to provide a comprehensive overview of the potential that audio-based inputs and multimodal approaches hold for LLMs.
@@
This work focuses on addressing these gaps:
  • Zero-shot (ZS) preprocessing and ZS inference for LLMs.
  • A setup of various tasks, including a multi-label classification task.
  • Comparison of prompts and their effects on model performance.
  • Analysis of results between audio and text modalities, their combination, and their evaluation through custom-formulated metrics (MSS and DRS).
  • Comparison of different models from various families and sizes.
  • Evaluation of the effects of few-shot (FS) learning on different tasks and models to assess its impact on performance.
In this paper, we will also address some limitations of the models, including how different prompts can impact their performance and accuracy. Variations in input phrasing may lead to differing results [9], which is an important consideration when evaluating LLMs. In the methodology 4 section, this approach for comparing models will be outlined, focusing on text-based and audio-based inputs. In the Experimental Setup 4.2, we will discuss the various tasks conducted to evaluate the models’ performance. Lastly, in the Results and Discussion 5 section, we will analyze and discuss how the models performed across different metrics, highlighting their strengths and weaknesses.
@@ 2 Related Work
To process these modalities, researchers employ techniques from machine learning (ML), deep learning (DL), and increasingly, large language models (LLMs). ML methods are often used for structured data and straightforward predictions, as detailed in Section 2.2.1, while DL techniques, like neural networks, can handle unstructured data such as raw audio or video, as detailed in Section 2.2.2. LLMs are particularly useful for analyzing text data, as they can understand and generate human language in a context-aware manner, making them highly effective for assessing linguistic and semantic features within mental health interviews, as detailed in Section2.2.3.
Combining modalities is common practice to enhance prediction accuracy and robustness. By integrating information from text, audio, and video, models can capture a broader spectrum of emotional and behavioral signals, leading to more reliable and nuanced assessments. This multimodal approach allows the model to compensate for the weaknesses of individual modalities and strengthen its ability to detect mental health conditions.
@@ 2.2 Processing Techniques
Once the textual, audio, and visual data are preprocessed, different Processing techniques are applied to predict mental health conditions like depression and PTSD. These models aim to leverage the cleaned and extracted features from each modality, learning patterns that can indicate the presence of psychological distress. The most commonly used approaches include Machine Learning (ML), Deep Learning (DL), and more recently, Large Language Models (LLMs), each of which contributes uniquely to the field.
@@ 2.2.2 Deep Learning (DL)
Additionally, David Gimeno-Gómez et al. (2024)[11] focused on a multimodal temporal model that processes non-verbal cues from various modalities, utilizing deep learning architectures to improve predictions in mental health detection. The combination of multiple inputs such as audio, visual, and text through deep learning models allows for more comprehensive analyses of participant behaviors.
2.2.3 Large Language Models (LLMs)
Large Language Models (LLMs) like BERT, GPT-3.5, and GPT-4 have become essential tools in analyzing text data, particularly when working with transcriptions from clinical interviews. LLMs excel at understanding the deeper semantic context and patterns within language, making them highly effective for predicting mental health conditions from text-based data.
@@ 2.2.3 Large Language Models (LLMs)
Clinton Lau et al. (2023)[24] applied prefix-tuning to large language models, such as GPT-4, to fine-tune their performance for specific tasks like depression severity estimation. By leveraging the contextual power of LLMs, they were able to capture subtle emotional cues from the patient transcripts that might be missed by traditional models.
@@ 3 Datasets
The Extended Distress Analysis Interview Corpus (E-DAIC) is an enhanced version of the DAIC-WOZ, designed to study psychological conditions such as anxiety, depression, and PTSD through semi-clinical interviews. The interviews are conducted by a human-controlled virtual agent named "Ellie" in a wizard-of-Oz (WoZ) setting or by an autonomous AI agent, both aiming to detect verbal and nonverbal indicators of mental illnesses. Developed as part of the DARPA Detection and Computational Analysis of Psychological Signals (DCAPS) program, this dataset is specifically crafted to advance the understanding and detection of psychological stress signals, with a particular focus on depression. The dataset is available through the University of Southern California’s Institute for Creative Technologies (USC ICT) and can be accessed by researchers through a data use agreement, ensuring ethical compliance and protection of participant data. Approval for the use of this dataset was obtained from USC ICT, emphasizing its adherence to institutional guidelines for studying psychological health conditions.
@@
  • Visual features including Pose, Gaze, and Action Units (AUs) summarized over set intervals.
  • Deep representations from CNN models like ResNet and VGG, and Densenet for spectral images converted from audio.
@@ 4.1 Evaluation Pipeline for Audio-Based Data
The proposed evaluation pipeline for audio-based data, as illustrated in Figure 2, begins with raw audio inputs. These audio samples can be directly provided to the model, or first transcribed into text using the Whisper Large-V3 model. In addition, the pipeline supports integrating both modalities—raw audio and transcribed text—together. After determining the preferred input format (audio only, text only, or a combination of both), a prompt engineering step is conducted. Here, carefully crafted task-specific prompts guide the model toward binary classification, severity classification, or multi-label classification tasks. These prompts are designed to ensure that the model receives clear instructions, aligned with the chosen input modality or modalities.
Once the input (audio, transcription, or both) is combined with the tailored prompts, the resulting prompt is passed to large language models (LLMs) for evaluation. This approach enables the assessment of the model’s zero-shot capabilities—evaluating how well it can perform classification tasks without prior fine-tuning or preprocessing. By examining LLM responses across different modalities and tasks, this pipeline provides insights into the model’s inherent ability to generalize, adapt, and accurately interpret a variety of input formats and instructions.
@@ 4.2 Experimental Setup
In this study, a comprehensive experimental setup was established to evaluate the effectiveness of Large Language Models in predicting mental health conditions, specifically depression and PTSD, using both text and audio modalities. The main objective of this experiment is to compare the performance of LLMs when processing textual data, such as transcriptions from interviews, against their performance in analyzing audio data that includes vocal features. By leveraging multimodal data, the aim is to assess how well each modality contributes to the accurate prediction of mental health conditions and whether combining them can enhance the overall predictive power of the models. In this setup, the audio modality provides information about vocal features, such as tone, pitch, and speech rate, which are indicative of emotional states. On the other hand, the text modality offers insights into the linguistic patterns and cognitive expressions of the participants, as derived from the transcriptions. Both data types are processed using state-of-the-art LLMs. Additionally, we explore the integration of both modalities to determine if this approach enhances model performance, providing a more comprehensive understanding of each participant’s mental state. This comparative study aims to provide insights into the strengths and limitations of each modality in mental health prediction and evaluate the benefits of combining them for more robust and accurate classification.
@@ 4.4 Transcription Process
One of the key features of Whisper’s architecture is its ability to handle multilingual transcription, background noise, and various accents with high precision. Whisper processes audio data in chunks, typically by converting the audio into spectrograms (a visual representation of sound) that can then be interpreted by the neural network. This process allows Whisper to extract meaningful patterns from the audio signal, even in challenging acoustic conditions, such as overlapping speech or background noise.
@@
"Have you ever served in the military? No. Have you ever been diagnosed with PTSD? Yes, I have. How long ago were you diagnosed? In, um, I don’t know. I was in the military. I was in the military. How long ago were you diagnosed? In February of 2011. What got you to seek help? I was attacked by a stalker and almost killed in November of 2009. He broke into my apartment and laid in wait for me and attacked me when I came in the door and tried to kill me. Do you still go to therapy now? I do."
@@ 4.5 LLMs Under Evaluation
Llama 3 70B[33] 70B Meta Groq Text
@@
In this study, we evaluated a diverse set of large language models (LLMs) to analyze their capabilities in handling both text and audio data. The models were selected based on factors such as parameter size, source, accessibility via APIs, and support for different data modalities. as well as those excluded, based on findings and recommendations from the referenced paper [9]. These criteria ensured a comprehensive comparison of models from various providers, including both proprietary and open-source options.
Table 4 summarizes the characteristics of the selected models, which vary significantly in terms of parameter size, ranging from lightweight models like Phi-3.5-mini (3.8 billion parameters) to much larger models such as Llama 3 70B. Additionally, the sources of the models encompass major technology companies like Google and Microsoft, as well as specialized AI firms such as Mistral AI and Meta. API accessibility is provided by multiple platforms, including Nvidia NIM, OpenAI, Groq, and Google’s Gemini API, enabling diverse deployment options for text and audio processing tasks.
For the audio analysis, Whisper was employed to transcribe audio files into text before inputting the results into text-focused models like Llama 3 and GPT-4o mini. In contrast, models supporting multimodal data, such as Gemini 1.5 Flash and Pro version 2, were directly fed audio data to evaluate their performance in handling both audio and text tasks. This approach allowed for a direct comparison of text-only versus multimodal model capabilities.
Several models in this evaluation, such as Phi-3.5-mini and Phi-3.5-MoE, were chosen due to their emerging relevance in multimodal and multilingual tasks. The inclusion of models with different quantization strategies, such as those used in Llama 3 70B and Mistral NeMo, highlights the trade-offs between model complexity and computational efficiency. Quantization, in some cases, helped optimize model performance, particularly in audio-capable models.
@@
Tabel 5 presents a summary of models that were excluded from the study because they did not meet the specific requirements needed for the analysis. The exclusion criteria were based on technical limitations that hindered the models’ ability to process the dataset effectively or misalignments between the models’ primary functionalities and the study’s objectives. Each model had distinct reasons for exclusion, ranging from input constraints and limitations in audio processing capabilities to being optimized for tasks that did not fit the study’s focus. By outlining these issues, the table helps clarify the rationale behind selecting alternative models that better match the study’s requirements for evaluating audio and text data.

### v2

@@ Abstract
Mental health disorders such as depression and PTSD are increasingly prevalent worldwide, creating an urgent need for innovative tools to support early diagnosis and intervention. This study investigates the potential of Large Language Models (LLMs) in multimodal diagnostics for mental health disorders, specifically depression and PTSD. Using the E-DAIC dataset, we evaluate performance across text and audio modalities, as well as their integration, using zero-shot inference without task-specific fine-tuning. The best performance for binary depression classification was achieved by Deepseek_R1, reaching a Balanced Accuracy (BA) of 79.8% and an F1 score of 0.75 using text input. For PTSD classification, GPT-4o mini attained a BA of 77% and an F1 score of 0.68. Significantly, models that support multimodal input consistently benefited from combining text and audio. For example, Gemini 2.5 Pro achieved an F1 score of 0.77 and BA of 77.8% when both modalities were integrated, representing a 3% improvement over text-only and 2.3% over audio-only inputs. This paper further investigates how modality fusion impacts models differently across binary, severity, multiclass, and multi-label tasks. It also introduces new evaluation metrics—Modal Superiority Score and Disagreement Resolution Score (DRS)—to assess model reliability and complementarity. These findings demonstrate the promise of LLMs, particularly in multimodal configurations, for enhancing clinical mental health assessments.
@@ 1 Introduction
In recent years, models like BERT[2] have advanced rapidly, demonstrating significant capabilities in tasks ranging from natural language processing to decision-making in specialized domains. [3] highlights the capabilities of LLMs as they are trained on vast and diverse corpora, enabling them to capture complex patterns in language. Their training allows them to generate coherent and contextually relevant text, making them highly adaptable across various applications. In healthcare, for example, LLMs can assist in diagnosing diseases, summarizing patient records, and even providing support for therapeutic interventions. Their ability to generalize across tasks has made them valuable in different fields, significantly enhancing efficiency and scalability.
LLMs show a remarkable ability to process and analyze vast amounts of text data, identifying and predicting psychiatric conditions by detecting patterns and subtle linguistic cues in patient communication. They excel at providing timely, scalable, and personalized assessments, aiding mental health professionals in diagnosis and treatment planning. LLMs provide scalable, efficient, and objective evaluations, enhancing diagnostic accuracy and personalization of treatment. However, they must be used carefully, as they may generate inaccurate responses or lead to over-reliance, requiring continuous monitoring to ensure safety and effectiveness[4][5].
Recent LLM advancements have enabled models to process not only text but also audio inputs directly  cite fang2024llamaomniseamlessspeechinteraction, providing valuable insights into mental health. Some models convert audio to text via a speech recognition system [6], analyzing the linguistic content for symptoms of mental disorders. However, newer models can directly analyze audio by detecting differences in tone, cadence, and speech patterns, which can reflect emotional states more accurately than text alone [7]. This approach allows for deeper insights into a patient’s mental health, potentially leading to more precise and early diagnosis of conditions like depression or anxiety.
In this paper, we compare two approaches for mental illness detection: text modality and audio modality, using different models to analyze each. Models processing text evaluate written transcripts to identify linguistic patterns, while models processing audio analyze direct raw speech to capture features such as tone and cadence. Additionally, we explore the integration of both modalities to determine if this enhances model performance. To quantitatively assess how the combined modality affects performance, we utilize our specially formulated metrics: the MSS and DRS. By evaluating both individual and integrated modalities, we aim to provide a comprehensive overview of the potential that audio-based inputs and multimodal approaches hold for LLMs.
@@
This study addresses those gaps by:
  • performing zero-shot (ZS) preprocessing and evaluation of multiple LLMs across diverse tasks, including multi-label classification;
  • systematically comparing distinct LLM families and examining how prompt design shapes their performance;
  • analyzing results for audio and text modalities—individually and in combination—using bespoke metrics (MSS and DRS);
  • investigating how few-shot (FS) learning affects model performance across the various tasks.
This paper also examines the models’ limitations, with particular attention to how prompt wording can influence their accuracy. Subtle changes in phrasing can lead to markedly different outputs [8], a crucial factor when evaluating LLMs. Section 4 details the comparison strategy for both text- and audio-based inputs, while Section 4.2 outlines the experimental tasks used to gauge performance. Section 5 then presents and interprets the findings, highlighting each model’s key strengths and weaknesses.
@@ 2 Related Work
To process these modalities, researchers employ techniques from machine learning (ML), deep learning (DL), and increasingly, LLMs. ML methods are often used for structured data and straightforward predictions, as detailed in Section 2.2.1. At the same time, DL techniques, like neural networks, can handle unstructured data such as raw audio or video, as detailed in Section 2.2.2. LLMs are particularly useful for analyzing text data, as they can understand and generate human language in a context-aware manner, making them highly effective for assessing linguistic and semantic features within mental health interviews, as detailed in Section 2.2.3.
Combining modalities is a common practice to enhance prediction accuracy and robustness. By integrating information from text, audio, and video, models can capture a broader spectrum of emotional and behavioral signals, leading to more reliable and nuanced assessments. This multimodal approach allows the model to compensate for the weaknesses of individual modalities and strengthen its ability to detect mental health conditions.
@@ 2.2 Processing Techniques
Once the textual, audio, and visual data are preprocessed, different Processing techniques are applied to predict mental health conditions like depression and PTSD. These models aim to leverage the cleaned and extracted features from each modality, learning patterns that can indicate the presence of psychological distress. The most commonly used approaches include Machine Learning (ML), Deep Learning (DL), and, more recently, LLMs, each of which contributes uniquely to the field.
@@ 2.2.2 Deep Learning (DL)
Additionally, David Gimeno-Gómez et al. (2024)[10] focused on a multimodal temporal model that processes non-verbal cues from various modalities, utilizing deep learning architectures to improve predictions in mental health detection. The combination of multiple inputs, such as audio, visual, and text, through deep learning models allows for more comprehensive analyses of participant behaviors.
2.2.3 Large Language Models
LLMs like BERT, GPT-3.5, and GPT-4 have become essential tools in analyzing text data, particularly when working with transcriptions from clinical interviews. LLMs excel at understanding the deeper semantic context and patterns within language, making them highly effective for predicting mental health conditions from text-based data.
@@ 2.2.3 Large Language Models
Clinton Lau et al. (2023)[23] applied prefix-tuning to large language models, such as GPT-4, to fine-tune their performance for specific tasks like depression severity estimation. By leveraging the contextual power of LLMs, they were able to capture subtle emotional cues from the patient transcripts that traditional models might miss.
@@ 3 Datasets
The Extended Distress Analysis Interview Corpus (E-DAIC) is an enhanced version of the DAIC-WOZ, designed to study psychological conditions such as anxiety, depression, and PTSD through semi-clinical interviews. The interviews are conducted by a human-controlled virtual agent named "Ellie" in a wizard-of-Oz (WoZ) setting or by an autonomous AI agent, both aiming to detect verbal and nonverbal indicators of mental illnesses. Developed as part of the DARPA Detection and Computational Analysis of Psychological Signals (DCAPS) program, this dataset is specifically crafted to advance the understanding and detection of psychological stress signals, with a particular focus on depression. The dataset is available through the University of Southern California’s Institute for Creative Technologies (USC ICT). It can be accessed by researchers through a data use agreement, ensuring ethical compliance and protection of participant data. Approval for the use of this dataset was obtained from USC ICT, emphasizing its adherence to institutional guidelines for studying psychological health conditions.
@@
  • Visual features including Pose, Gaze, and Action Units (AUs), summarized over set intervals.
  • Deep representations from CNN models like ResNet, VGG, and Densenet for spectral images converted from audio.
@@ 4.1 Evaluation Pipeline for Audio-Based Data
The proposed evaluation pipeline for audio-based data, as illustrated in Figure 2, begins with raw audio inputs. These audio samples can be directly provided to the model or first transcribed into text using the Whisper Large-V3 model. In addition, the pipeline supports integrating both modalities—raw audio and transcribed text. After determining the preferred input format (audio only, text only, or a combination of both), a prompt engineering step is conducted. Here, carefully crafted task-specific prompts guide the model toward binary classification, severity classification, or multi-label classification tasks. These prompts are designed to ensure that the model receives clear instructions, aligned with the chosen input modality or modalities.
Once the input (audio, transcription, or both) is combined with the tailored prompts, the resulting prompt is passed to LLMs for evaluation. This approach enables the assessment of the model’s zero-shot capabilities, evaluating how well it can perform classification tasks without prior fine-tuning or preprocessing. By examining LLM responses across different modalities and functions, this pipeline provides insights into the model’s inherent ability to generalize, adapt, and accurately interpret a variety of input formats and instructions.
@@ 4.2 Experimental Setup
In this study, a comprehensive experimental setup was established to evaluate the effectiveness of LLMs in predicting mental health conditions, specifically depression and PTSD, using both text and audio modalities. The main objective of this experiment is to compare the performance of LLMs when processing textual data, such as transcriptions from interviews, against their performance in analyzing audio data that includes vocal features. By leveraging multimodal data, the aim is to assess how well each modality contributes to the accurate prediction of mental health conditions and whether combining them can enhance the overall predictive power of the models. In this setup, the audio modality provides information about vocal features, such as tone, pitch, and speech rate, which are indicative of emotional states. On the other hand, the text modality offers insights into the linguistic patterns and cognitive expressions of the participants, as derived from the transcriptions. Both data types are processed using state-of-the-art LLMs. Additionally, we explore the integration of both modalities to determine if this approach enhances model performance, providing a more comprehensive understanding of each participant’s mental state. This comparative study aims to provide insights into the strengths and limitations of each modality in mental health prediction and evaluate the benefits of combining them for more robust and accurate classification.
@@ 4.4 Transcription Process
One of the key features of Whisper’s architecture is its ability to handle multilingual transcription, background noise, and various accents with high precision. Whisper processes audio data in chunks, typically by converting the audio into spectrograms (a visual representation of sound) that the neural network can then interpret. This process allows Whisper to extract meaningful patterns from the audio signal, even in challenging acoustic conditions, such as overlapping speech or background noise.
@@
"Have you ever served in the military? No. Have you ever been diagnosed with PTSD? Yes, I have. How long ago were you diagnosed? In, um, I don’t know. I was in the military. I was in the military. How long ago were you diagnosed? In February of 2011. What got you to seek help? A stalker attacked me and almost killed me in November of 2009. He broke into my apartment and lay in wait for me, and attacked me when I came in the door, and tried to kill me. Do you still go to therapy now? I do."
@@ 4.5 LLMs Under Evaluation
Llama 3.1 70B [32] 70B Meta Groq Text
@@
Deepseek_R1 [37] Proprietary Deepseek Deepseek API Text
Llama 3.3 70B [32] 70B Groq Groq API Text
@@
Gemini 2.5 Pro [38] Proprietary Google Google’s Gemini API Text & Audio
Gemini 2 Flash [38] Proprietary Google Google’s Gemini API Text & Audio
Gemini Flash Thinking [38] Proprietary Google Google’s Gemini API Text & Audio
Gemini 2 Flash Lite [38] Proprietary Google Google’s Gemini API Text & Audio
In this study, we evaluated a diverse set of large language models (LLMs) to analyze their capabilities in handling both text and audio data. The models were selected based on factors such as parameter size, source, accessibility via APIs, and support for different data modalities. Additionally, models were excluded based on findings and recommendations from the referenced paper [8]. These criteria ensured a comprehensive comparison of models from various providers, including both proprietary and open-source options.
Table 4 Table 4 summarizes the characteristics of the selected models, which vary significantly in terms of parameter size, from lightweight models like Phi-3.5-mini (3.8B parameters) to larger ones such as Llama 3.1 and 3.3 (both 70B). The sources include major companies like Google, Microsoft, and Meta, as well as specialized firms like Mistral AI and Deepseek. API accessibility is provided through a range of platforms, including Nvidia NIM, OpenAI API, Groq, Azure AI API, and Google’s Gemini API, allowing flexible deployment options across text and audio modalities.
For the audio analysis, Whisper was employed to transcribe audio files into text before inputting the results into text-focused models like Llama 3 and GPT-4o mini. In contrast, models supporting multimodal data, such as Gemini 1.5 Flash, Gemini 1.5 Pro, and Gemini 2.5 Pro, were directly fed audio data to evaluate their performance in handling both audio and text tasks. This approach allowed for a direct comparison of text-only versus multimodal model capabilities.
Several models in this evaluation, such as Phi-3.5-mini and Phi-3.5-MoE, were chosen due to their emerging relevance in multimodal and multilingual tasks. The inclusion of models with different quantization strategies, such as those used in Llama 3.1 70B and Mistral NeMo, highlights the trade-offs between model complexity and computational efficiency. Quantization, in some cases, helped optimize model performance, particularly in audio-capable models.
@@
Table 5 presents a summary of models that were excluded from the study because they did not meet the specific requirements needed for the analysis. The exclusion criteria were based on technical limitations that hindered the models’ ability to process the dataset effectively or misalignments between the models’ primary functionalities and the study’s objectives. Each model had distinct reasons for exclusion, ranging from input constraints and limitations in audio processing capabilities to being optimized for tasks that did not fit the study’s focus. By outlining these issues, the table helps clarify the rationale behind selecting alternative models that better match the study’s requirements for evaluating audio and text data.